Susan runs 3 miles in 22 minutes. At the same rate, how many miles would she run in 77 minutes?
Found 2 solutions by lynnlo, jpvn2015:
Answer by lynnlo(4176)

3/22=x/77======make your equation
22x=231========multiply
x=10.5=========divide both sides by 22 to get what X means
X=10.5

Answer by jpvn2015(54)

Set this problem into cross multiplication format:
3/22 = y/77
After the multiplication: 22y=77(3) = 231
Thus, 22y=231
Divide both sides by 22: y=10.5
Therefore, 10.5 miles is the answer.